OpenStudy (firejay5):

I need help figuring this out of the geometric mean of this word problem. I'll draw the figure out so be patient. Medal will be rewarded for your patronage! :D

OpenStudy (firejay5):



OpenStudy (anonymous):

 <i+<2+(<5+<6)=180 or <1+<2+90=180,<1+<2=180-90=90 ...(1) <1+<5+90=180, <1+<5=90 .....(2) <2+<6+90=180, <2+<6=90 ......(3) from (1) and (2), <1+<2=<1+<5 or <2=<5 from (1) and (3) <1+<2=<2+<6 <1=<6 triangles (2) and (3) are similar. Hence sides are proportional. \[\frac{ BD }{ DC }=\frac{ AD }{ BD },(BD)^{2}=AD*DC=x*9=9x,BD=3\sqrt{x}\]